# The Basic Reproduction Number for Petri Net Models: A Next-Generation Matrix Approach

### v2

@@ 4.4 Basic COVID Model
For $V$, the transition matrix, we start by finding $\mathcal{V}$ based on the arc weights between infected places and the arc weights leaving the infected places as shown in Equation 3. Thus
$\mathcal{V}_{1,2,3,4}(x)$ $=$ $\begin{bmatrix}\sigma E&0&0&0\\ -r\sigma E&\gamma_{a}I_{a}&0&0\\ -(1-r)\sigma E&0&\varphi_{S}I_{S}+\gamma_{S}I_{S}+\delta_{S}I_{S}&0\\ 0&0&-\varphi_{S}I_{S}&\gamma_{h}I_{H}\end{bmatrix}$
Then taking the partial derivative with respect $E,I_{a},I_{S},I_{h}$ for the first, second, third, and fourth columns, respectively, gives us
$V$ $=$ $\begin{bmatrix}\sigma&0&0&0\\ -r\sigma&\gamma_{a}&0&0\\ -(1-r)\sigma&0&\varphi_{S}+\gamma_{S}+\delta_{S}&0\\ 0&0&\varphi_{S}&\gamma_{h}\end{bmatrix}$
@@
$R_{0}=\varrho(FV^{-1})=\left(\left(\frac{\beta_{a}r}{\gamma_{a}}\right)+\left(\frac{\beta_{s}(1-r)}{\varphi_{s}+\gamma_{s}+\delta_{s}}\right)+\left(\frac{\beta_{h}(1-r)\varphi_{s}}{(\varphi_{s}+\gamma_{s}+\delta_{s})(\gamma_{h}+\delta_{h})}\right)\right)\left(\frac{S^{*}}{N^{*}}\right).$
The values for the PN $F$ and $V$ are equivalent to that values for the ODE $F$ and $V$ and the resulting $R_{0}$ matches the $R_{0}$ for the ODE system in the Gumel paper [21] seen in Equations (2) and (3).
@@ 4.5 Nonlinear System
The paper by Rohith et al. [22] involves nonlinear compartment-to-compartment dynamics. These types of dynamics are more applicable when wanting to involve the dynamics of social aspects of a disease spread, drug pharmacokinetics, or disease cell spread within an individual. The analysis of the nonlinear system underscores the robustness of the NGMPN framework, demonstrating its applicability even when departing from standard mass-action assumptions to incorporate more realistic, behavior-driven transmission dynamics. The parameters $\mu,\beta,\alpha,\sigma,$ and $\gamma$ represent the birth/death rate, the per capita contact rate, the psychological or inhibitory effect, rate of infectiousness, and the recovery rate, respectively.
$\frac{dS}{dt}$ $=$ $\mu-\frac{\beta SI}{1+\alpha I^{2}}-\mu S$ (24)
$\frac{dE}{dt}$ $=$ $\frac{\beta SI}{1+\alpha I^{2}}-\sigma E-\mu E$ (25)
$\frac{dI}{dt}$ $=$ $\sigma E-\gamma I-\mu I$ (26)
$\frac{dR}{dt}$ $=$ $\gamma I-\mu R$ (27)
Figure 6: Non-linear SEIR Variable Arc Weight PN Model
From Figure 6 we can see that there are two infected places $(E,I)$. Thus our resulting $F$ and $V$ matrices will be two by two. Then, from inspecting the PN model, we see that $\frac{\beta SI}{1+\alpha I^{2}}$ is the only way that the new infected population enters. Thus the resulting $\mathcal{F}$ is
@@
Then finding the DFE of the PN system to be $(S^{*},E^{*},I^{*},R^{*})=(1,0,0,0)$. We then take the partial derivative of each element of $\mathcal{F}_{1,2}$ with respect to $E,I$ for the first and second columns, respectively, to give the transmission matrix $F$,
@@
For the transition matrix $V$, we start by finding $\mathcal{V}$ based on the arc weights between infected places and the arc weights leaving the infected places yielding
@@
which is the same $R_{0}$ value found in the Rohith paper [22].
@@ 4.6 Patch System Overview
The paper Multi-patch and multi-group epidemic models: a new framework by Bichara and Iggidr [23] presents a multi-patch and multi-group model for infectious disease transmission. The model accommodates interactions occurring across an arbitrary number of patches and groups where the infection propagates. The compartments $S_{i},E_{i},I_{i},R_{i}$ still represent susceptible, exposed, infected, and recovered, respectively, just now for patch $i$. Similarly, the parameters $\Pi,\beta,\mu,\eta,\upsilon,\gamma,$ and $\delta$ represent birth/immigration, rate of infection, natural death rate, loss of immunity rate, rate of becoming infectious, rate of recovery and rate of disease related death, respectively. The parameter $\beta$ is a patch specific term while $\Pi,\mu,\eta,\upsilon,\gamma,$ and $\delta$ are group specific terms. People of group $i$ spend on average some time in patch $j$,$j=1,...,v$. The susceptible, exposed, infected, and recovered populations of group $i$ spend $m_{ij},n_{ij},p_{ij},$ and $q_{ij}$ proportion of time in patch $j$ for $j=1,...,v$. , , and proportion of times respectively in Patch j, for . The ODEs are outlined here
$\frac{dS_{i}}{dt}$ $=$ $\Pi_{i}-\sum_{j=1}^{v}\beta_{j}m_{ij}S_{i}\frac{\sum_{k=1}^{u}p_{kj}I_{k}}{\sum_{k=1}^{u}(m_{kj}S_{k}+n_{kj}E_{k}+p_{kj}I_{k}+q_{kj}R_{k})}-\mu_{i}S_{i}+\eta_{i}R_{i}$ (28)
$\frac{dE_{i}}{dt}$ $=$ $\sum_{j=1}^{v}\beta_{j}m_{ij}S_{i}\frac{\sum_{k=1}^{u}p_{kj}I_{k}}{\sum_{k=1}^{u}(m_{kj}S_{k}+n_{kj}E_{k}+p_{kj}I_{k}+q_{kj}R_{k})}-\upsilon_{i}E_{I}-\mu_{i}E_{i}$ (29)
$\frac{dI_{i}}{dt}$ $=$ $\upsilon_{i}E_{I}-\gamma_{i}I_{i}-\delta_{i}I_{i}-\mu_{i}I_{i}$ (30)
@@
4.6.1 Two Patch System
We take the ODEs from Bichara and Iggidr [23] Equations 28-31 and lay them out explicitly for two patches and two groups. Though the ODE and corresponding PN model only are for this small patch and group size, this model could be extended to any size with the interactions between any two given patches being equivalent to the system laid out below. The multi-patch model highlights the scalability of our approach to spatially structured populations, providing a tool to investigate the impact of geographic factors and population mobility on disease spread.
$\frac{dS_{1}}{dt}$ $=$ $\Pi_{1}-\left(\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}+\right....$ (32)
$\beta_{2}m_{12}S_{1}\left.\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)...$
$-\mu_{1}S_{1}+\eta_{1}R_{1}$
$\frac{dS_{2}}{dt}$ $=$ $\Pi_{2}-\left(\beta_{1}m_{21}S_{2}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}+\right....$ (33)
${}_{22}\beta_{2}m_{22}S_{2}\left.\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)...$
$-\mu_{2}S_{2}+\eta_{2}R_{2}$
$\frac{dE_{1}}{dt}$ $=$ $\left(\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}+\right....$ (34)
$\beta_{2}m_{12}S_{1}\left.\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{k}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)...$
$-\upsilon_{1}E_{1}-\mu_{1}E_{1}$
$\frac{dE_{2}}{dt}$ $=$ $\left(\beta_{1}m_{21}S_{2}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}+\right....$ (35)
$\beta_{2}m_{22}S_{2}\left.\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{k}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)...$
$-\upsilon_{2}E_{2}-\mu_{2}E_{2}$
$\frac{dI_{1}}{dt}$ $=$ $\upsilon_{1}E_{1}-\gamma_{1}I_{1}-\delta_{1}I_{1}-\mu_{1}I_{1}$ (36)
$\frac{dI_{2}}{dt}$ $=$ $\upsilon_{2}E_{2}-\gamma_{2}I_{2}-\delta_{2}I_{2}-\mu_{2}I_{2}$ (37)
$\frac{dR_{1}}{dt}$ $=$ $\gamma_{1}I_{1}-\eta_{1}R_{1}-\mu_{1}R_{1}$ (38)
$\frac{dR_{2}}{dt}$ $=$ $\gamma_{2}I_{2}-\eta_{2}R_{2}-\mu_{i}R_{2}$ (39)
Figure 7: Patch Variable Arc Weight PN Model
To simplify, we define $\Theta_{1}=\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}$,
@@ 4.6.1 Two Patch System
Then we obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the non-infected places token values $(S_{1}^{*},S_{2}^{*},E_{1}^{*},E_{2}^{*},I_{1}^{*},I_{2}^{*},R_{1}^{*},R_{2}^{*})=(\frac{\Pi_{1}}{\mu_{1}},\frac{\Pi_{2}}{\mu_{2}},0,0,0,0,0,0)$. Then the matrix $\mathcal{F}$ by looking at the arc weights of arcs going into the infected compartments giving us
@@
From this we find the rate of change with respect to $E_{1}$ in the first column, $E_{2}$ in the second column, $I$ in the third column, and evaluate at the DFE to give
$F$ $=$ $\begin{bmatrix}0&0&\frac{\beta_{1}m_{11}p_{11}S_{1}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{12}p_{12}S_{1}}{m_{12}S_{1}+m_{22}S_{2}}&\frac{\beta_{1}m_{11}p_{12}S_{1}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{12}p_{22}S_{1}}{m_{12}S_{1}+m_{22}S_{2}}\\ 0&0&\frac{\beta_{1}m_{21}p_{11}S_{2}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{22}p_{12}S_{2}}{m_{12}S_{1}+m_{22}S_{2}}&\frac{\beta_{1}m_{21}p_{21}S_{2}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{22}p_{22}S_{2}}{m_{12}S_{1}+m_{22}S_{2}}\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}.$
@@
For $\mathcal{V}$, we look at arc weights of arcs leaving the infected places and transitioning from one infected place to another yielding
$\mathcal{V}_{1,2,3,4}(x)$ $=$ $\begin{bmatrix}\nu_{1}E_{1}+\mu_{1}E_{1}&0&0&0\\ 0&\nu_{2}E_{2}+\mu_{2}E_{2}&0&0\\ -\nu_{1}E_{1}&0&\gamma_{1}I_{1}+\delta_{1}I_{1}+\mu_{1}I_{1}&0\\ 0&-\nu_{2}E_{2}&0&\gamma_{2}I_{2}+\delta_{2}I_{2}+\mu_{2}I_{2}\end{bmatrix}.$
$V$ $=$ $\begin{bmatrix}\nu_{1}+\mu_{1}&0&0&0\\ 0&\nu_{2}+\mu_{2}&0&0\\ -\nu_{1}&0&\gamma_{1}+\delta_{1}+\mu_{1}&0\\ 0&-\nu_{2}&0&\gamma_{2}+\delta_{2}+\mu_{2}\end{bmatrix}$
$V^{-1}$ $=$ $\begin{bmatrix}\frac{1}{\nu_{1}+\mu_{1}}&0&0&0\\ 0&\frac{1}{\nu_{2}+\mu_{2}}&0&0\\ \frac{\nu_{1}}{(\nu_{1}+\mu_{1})(\gamma_{1}+\delta_{1}+\mu_{1})}&0&\frac{1}{\gamma_{1}+\delta_{1}+\mu_{1}}&0\\ 0&\frac{\nu_{2}}{(\nu_{2}+\mu_{2})(\gamma_{2}+\delta_{2}+\mu_{2})}&0&\frac{1}{\gamma_{2}+\delta_{2}+\mu_{2}}\end{bmatrix}.$
@@
Now that we have $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and determining the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}F\cdot V^{-1}\end{bmatrix}=$

### v3

@@ 4.5 Basic COVID-19 Model
For $V$, the transition matrix, we start by finding $\mathcal{V}$ based on the arc weights between infected places and the arc weights leaving the infected places as shown in Equation (3). Thus,
$\mathcal{V}_{1,2,3,4}(x)$ $=\begin{bmatrix}\sigma E&0&0&0\\ -r\sigma E&\gamma_{a}I_{a}&0&0\\ -(1-r)\sigma E&0&\varphi_{S}I_{S}+\gamma_{S}I_{S}+\delta_{S}I_{S}&0\\ 0&0&-\varphi_{S}I_{S}&\gamma_{h}I_{h}+\delta_{h}I_{h}\end{bmatrix}$.
Then, taking the partial derivative with respect to $E,I_{a},I_{S},I_{h}$ for the first, second, third, and fourth columns, respectively, gives us
$V$ $=\begin{bmatrix}\sigma&0&0&0\\ -r\sigma&\gamma_{a}&0&0\\ -(1-r)\sigma&0&\varphi_{S}+\gamma_{S}+\delta_{S}&0\\ 0&0&\varphi_{S}&\gamma_{h}+\delta_{h}\end{bmatrix}$.
@@
$R_{0}=\varphi(FV^{-1})=\left(\left(\frac{\beta_{a}r}{\gamma_{a}}\right)+\left(\frac{\beta_{s}(1-r)}{\varphi_{s}+\gamma_{s}+\delta_{s}}\right)+\left(\frac{\beta_{h}(1-r)\varphi_{s}}{(\varphi_{s}+\gamma_{s}+\delta_{s})(\gamma_{h}+\delta_{h})}\right)\right)\left(\frac{S^{*}}{N^{*}}\right).$
The values for the Petri Net $F$ and $V$ are equivalent to those values for the ODE $F$ and $V$, and the resulting $R_{0}$ matches the $R_{0}$ for the ODE system in the Gumel paper Gumel et al. (2021) seen in Equations (2) and (3).
@@ 4.6 Nonlinear System
The paper by Rohith et al. Rohith and Devika (2020) involves nonlinear compartment-to-compartment dynamics. These types of dynamics are more applicable when considering the social aspects of disease spread, drug pharmacokinetics, or disease cell spread within an individual. The analysis of the nonlinear system underscores the robustness of the NGMPN framework, demonstrating its applicability even when departing from standard mass-action assumptions to incorporate more realistic, behavior-driven transmission dynamics. The parameters $\mu,\beta,\alpha,\sigma,$ and $\gamma$ represent the birth/death rate, the per capita contact rate, the psychological or inhibitory effect, the rate of infectiousness, and the recovery rate, respectively. The nonlinear system is described by the following system of equations:
$\frac{dS}{dt}$ $=\mu-\frac{\beta SI}{1+\alpha I^{2}}-\mu S,$ (27)
$\frac{dE}{dt}$ $=\frac{\beta SI}{1+\alpha I^{2}}-\sigma E-\mu E,$ (28)
$\frac{dI}{dt}$ $=\sigma E-\gamma I-\mu I,$ (29)
$\frac{dR}{dt}$ $=\gamma I-\mu R.$ (30)
From Figure 7 we can see that there are two infected places $(E,I)$. Thus, our resulting $F$ and $V$ matrices will be two by two. Then, from inspecting the PN model, we see that $\frac{\beta SI}{1+\alpha I^{2}}$ is the only way that the new infected population enters. Thus the resulting $\mathcal{F}$ is
@@
Figure 7: A variable arc weight nonlinear SEIR PN mapped from Equations (27)–(30). The arc weights used in forming matrix $F_{i}(x)$ are outlined in red, and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
Then, finding the DFE of the PN system to be $(S^{*},E^{*},I^{*},R^{*})=(1,0,0,0)$. We then take the partial derivative of each element of $\mathcal{F}_{1,2}$ with respect to $E,I$ for the first and second columns, respectively, to give the transmission matrix $F$,
@@
For the transition matrix $V$, we start by finding $\mathcal{V}$ based on the arc weights between infected places and the arc weights leaving the infected places, yielding
@@
which is the same $R_{0}$ value found in the Rohith paper Rohith and Devika (2020). This nonlinear example, with its functional transmission term of $\frac{\beta SI}{1+\alpha I^{2}}$ represented by a single arc in Figure 7, perfectly illustrates the strength of our framework in handling VAPN models. A geometric, path-summing method based on constant-rate transitions would be difficult to apply to such a compact representation.
@@ 4.7 Patch System Overview
[…]
$\frac{dS_{i}}{dt}$ $=\Pi_{i}-\sum_{j=1}^{n}\beta_{j}m_{ij}S_{i}\frac{\sum_{k=1}^{u}p_{kj}I_{k}}{\sum_{k=1}^{u}(m_{kj}S_{k}+n_{kj}E_{k}+p_{kj}I_{k}+q_{kj}R_{k})}-\mu_{i}S_{i}+\eta_{i}R_{i},$ (31)
$\frac{dE_{i}}{dt}$ $=\sum_{j=1}^{n}\beta_{j}m_{ij}S_{i}\frac{\sum_{k=1}^{u}p_{kj}I_{k}}{\sum_{k=1}^{u}(m_{kj}S_{k}+n_{kj}E_{k}+p_{kj}I_{k}+q_{kj}R_{k})}-\upsilon_{i}E_{I}-\mu_{i}E_{i},$ (32)
$\frac{dI_{i}}{dt}$ $=\upsilon_{i}E_{I}-\gamma_{i}I_{i}-\delta_{i}I_{i}-\mu_{i}I_{i},$ (33)
@@
Two Patch System
We take the ODEs from Bichara and Iggidr Bichara and Iggidr (2018) Equations (31)–(34) and lay them out explicitly for two patches and two groups. Although the ODE and corresponding PN model are only applicable to this small patch and group size, this model can be extended to any size, with the interactions between any two given patches being equivalent to the system laid out below. The multi-patch model highlights the scalability of our approach to spatially structured populations, providing a tool to investigate the impact of geographic factors and population mobility on disease spread. The Two Patch system is described by the following system of equations:
$\frac{dS_{1}}{dt}$ $=\Pi_{1}-\left(\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}\right$.+
$\qquad\left.\beta_{2}m_{12}S_{1}\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)-\mu_{1}S_{1}+\eta_{1}R_{1},$ (35)
$\frac{dS_{2}}{dt}$ $=\Pi_{2}-\left(\beta_{1}m_{21}S_{2}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}\right$.+
$\qquad\left.\beta_{2}m_{22}S_{2}\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)-\mu_{2}S_{2}+\eta_{2}R_{2},$ (36)
$\frac{dE_{1}}{dt}$ $=\left(\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}\right$.+
$\qquad\left.\beta_{2}m_{12}S_{1}\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)-\upsilon_{1}E_{1}-\mu_{1}E_{1},$ (37)
$\frac{dE_{2}}{dt}$ $=\left(\beta_{1}m_{21}S_{2}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}\right$.+
$\qquad\left.\beta_{2}m_{22}S_{2}\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)-\upsilon_{2}E_{2}-\mu_{2}E_{2},$ (38)
$\frac{dI_{1}}{dt}$ $=\upsilon_{1}E_{1}-\gamma_{1}I_{1}-\delta_{1}I_{1}-\mu_{1}I_{1},$ (39)
$\frac{dI_{2}}{dt}$ $=\upsilon_{2}E_{2}-\gamma_{2}I_{2}-\delta_{2}I_{2}-\mu_{2}I_{2},$ (40)
$\frac{dR_{1}}{dt}$ $=\gamma_{1}I_{1}-\eta_{1}R_{1}-\mu_{1}R_{1},$ (41)
$\frac{dR_{2}}{dt}$ $=\gamma_{2}I_{2}-\eta_{2}R_{2}-\mu_{2}R_{2}.$ (42)
From Figure 8 we find there are two infected places ($E$, II) per patch. Thus, with two patches, our resulting $F$ and $V$ matrices will be four by four.
Figure 8: A variable arc weight two patch SEIR PN mapped from Equations (35)–(42). The arc weights used in forming matrix $F_{i}(x)$ are outlined in red, and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
To simplify, we define
$\small\Theta_{1}=\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})},$
@@
Then we obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the non-infected places token values $(S_{1}^{*},S_{2}^{*},E_{1}^{*},E_{2}^{*},I_{1}^{*},I_{2}^{*},R_{1}^{*},R_{2}^{*})=(\frac{\Pi_{1}}{\mu_{1}},\frac{\Pi_{2}}{\mu_{2}},0,0,0,0,0,0)$. Then the matrix $\mathcal{F}$ by looking at the arc weights of arcs going into the infected compartments, giving us
@@
Finding the rate of change with respect to the infected states and evaluating at the DFE yields
$F$ $=\begin{bmatrix}0&0&\frac{\beta_{1}m_{11}p_{11}S_{1}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{12}p_{12}S_{1}}{m_{12}S_{1}+m_{22}S_{2}}&\frac{\beta_{1}m_{11}p_{12}S_{1}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{12}p_{22}S_{1}}{m_{12}S_{1}+m_{22}S_{2}}\vskip 8.0pt\\ 0&0&\frac{\beta_{1}m_{21}p_{11}S_{2}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{22}p_{12}S_{2}}{m_{12}S_{1}+m_{22}S_{2}}&\frac{\beta_{1}m_{21}p_{21}S_{2}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{22}p_{22}S_{2}}{m_{12}S_{1}+m_{22}S_{2}}\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$.
@@
For $\mathcal{V}$, we find arcs leaving the infected places and transitioning between infected places, yielding
$\mathcal{V}_{1,2,3,4}(x)$ $=\begin{bmatrix}\upsilon_{1}E_{1}+\mu_{1}E_{1}&0&0&0\\ 0&\upsilon_{2}E_{2}+\mu_{2}E_{2}&0&0\\ -\upsilon_{1}E_{1}&0&\gamma_{1}I_{1}+\delta_{1}I_{1}+\mu_{1}I_{1}&0\\ 0&-\upsilon_{2}E_{2}&0&\gamma_{2}I_{2}+\delta_{2}I_{2}+\mu_{2}I_{2}\end{bmatrix}$.
We then find the rate of change with respect to the infected place $I$, and evaluate at the DFE. From this, we find the inverse to give
$V$ $=\begin{bmatrix}\upsilon_{1}+\mu_{1}&0&0&0\\ 0&\upsilon_{2}+\mu_{2}&0&0\\ -\upsilon_{1}&0&\gamma_{1}+\delta_{1}+\mu_{1}&0\\ 0&-\upsilon_{2}&0&\gamma_{2}+\delta_{2}+\mu_{2}\end{bmatrix}$,
$V^{-1}$ $=\begin{bmatrix}\frac{1}{\upsilon_{1}+\mu_{1}}&0&0&0\\ 0&\frac{1}{\upsilon_{2}+\mu_{2}}&0&0\\ \frac{\upsilon_{1}}{(\upsilon_{1}+\mu_{1})(\gamma_{1}+\delta_{1}+\mu_{1})}&0&\frac{1}{\gamma_{1}+\delta_{1}+\mu_{1}}&0\\ 0&\frac{\upsilon_{2}}{(\upsilon_{2}+\mu_{2})(\gamma_{2}+\delta_{2}+\mu_{2})}&0&\frac{1}{\gamma_{2}+\delta_{2}+\mu_{2}}\end{bmatrix}$.
@@
Now that we have $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and determining the dominant eigenvalue. Thus,